OpenStudy (anonymous):

Please help! Find the excluded value of the rational expression. (2x + 6)/(4x – 8) A. -3 B. -2 C. 0 D. 2 A, B, C, or D? Please explain (:

OpenStudy (jusaquikie):

instead of x put in a number from your answers and see what you get

OpenStudy (jusaquikie):

(2(-3)+6) / (4(-3)-8) = 0/-20 (2(-2)+6) / (4(-2)-8) =2/-16 (2(0)+6) / (4(0)-8) =0/0 (2(2)+6) / (4(2)-8) =10/0

OpenStudy (jusaquikie):

that didn't really get us anywhere let's just use algebra \[\frac{ (2x+6) }{(4x-8) } = \frac{ 2(x+3) }{ 4(x-2) }\]

OpenStudy (jusaquikie):

we want to find the value where the expression is undefined which happens when the denominator is zero. so what we have to do is make the denominator zero so what would make \[4(x-2) =0?\] when x=2 because \[4(2-2)=4(0)=0\] so D is your answer

OpenStudy (anonymous):

Thank You!